# Gauge potentials and vortices in the Fock space of a pair of periodically driven Bose-Einstein condensates

### v1

@@ Abstract
We perform a theoretical study of the coupled dynamics of two species of Bose-Einstein condensates (BECs) in a double well potential where both the tunneling and the interatomic interactions are driven periodically in time. The population difference between the wells of each species gives rise to a two dimensional lattice in Fock space with dimensions given by the number of atoms in each BEC. We use a Floquet analysis to derive an effective Hamiltonian that acts in this Fock space and find that it contains an artificial gauge field. This system simulates noninteracting particles in a tight binding lattice subject to an additional harmonic potential and vector potential. When the intra-species interactions are attractive there is a critical value at which the ground state of the Floquet operator undergoes a transition from a Gaussian state to a quantized vortex state in Fock space. The transition can be quantified in terms of the angular momentum as well as the entanglement entropy of the ground state with both showing sudden jumps as the intra-species interactions become stronger. The stability of the vortex state vanishes in the thermodynamic limit.
@@ I Introduction
Another useful tool for quantum simulation is provided by artificial gauge potentials. These mimic the action of electric and magnetic fields on charges irrespective of whether the underlying physical particles are charged or not [20, 21, 22, 23]. For ultracold atoms, a popular experimental configuration is to load them into an optical lattice which can then be periodically modulated to generate effective gauge potentials [24, 25, 26, 27, 28, 29, 30, 31, 32]. The ability to simulate gauge fields in a controlled and tuneable manner has opened up new possibilities for investigating topological phases [33, 34, 35, 36], exotic phases of matter [38, 39], and the study of quantum Hall physics [11, 40] on entirely different physical platforms to their original settings. Artificial gauge fields not only provide a versatile experimental method for probing fundamental aspects of quantum mechanics and condensed matter physics but also enable new technological applications in quantum information processing and quantum simulation [41, 42, 43].
@@
Fock states provide yet another possible set of states that can be used as building blocks for synthetic dimensions. Unlike the more traditionally used internal and external states listed above, Fock states are defined via the occupation numbers of particles in the different available modes leading to the concept of a Fock-state lattice (FSL) [45, 44]. Systems with a large number of particles and a limited number of modes are particularly well-suited for constructing large synthetic dimensions in this way because each mode acts as a single synthetic dimension, and the number of particles that can occupy a mode serves as the length of the dimension. In this context, particles with bosonic statistics are preferable due to their ability to occupy the same mode simultaneously. One example of a FSL is provided by photon states in optical cavities; when two or three optical cavities are coupled via a two-level atom such a FSL can realize the Su-Schrieffer-Heeger (SSH) model or a Lifshitz topological phase transition between a semimetal and a three band insulator, respectively [45, 46]. Likewise, experiments with superconducting circuits have successfully constructed one- and two-dimensional FSLs which can mimic both the SSH and Haldane models [47]. Of particular relevance to the current paper is a recent proposal by one of us (JM) based on two species of atomic Bose-Einstein condensates (BECs) confined in a double well potential which realizes a two-dimensional FSL [48]. This system can simulate strong synthetic gauge fields by intense periodic driving of the interactions between the two species of atoms leading to topological behavior in the form of chiral edge states.
Ever since the pioneering work by Onsager [49] and Feynman [50], quantized vortices have been regarded as the quintessential example of topological excitations. The original system of interest was liquid helium II, and subsequent experiments by Hall and Vinen [51, 52] confirmed that the circulation $\oint\mathbf{v}_{s}.d\mathbf{r}$ is quantized in units of $h/m$, where $h$ is Planck’s constant, $m$ is the mass of a helium atom. $\mathbf{v}_{s}$ is the velocity of the superfluid component. Shortly afterwards it was realized that quantized vortices also form in type II superconductors in the presence of magnetic fields [53, 54]. In this latter case the circulating current is charged and each vortex encloses a quantum of flux $h/2e$ where $e$ is the electron charge. More recently, quantized vortices have been intensively studied in neutral atomic BECs where an external trapping potential such as an optical lattice and/or harmonic trap often plays an important role [55, 56, 57]. Quantized vortices can also be generated in other types of fields such as optical beams [58, 59], and these find technological applications, e.g. in superresolution fluorescence microscopy [60].
@@
In this paper we study vortices in a FSL by combining elements of all the above mentioned systems. Specifically, we perform a theoretical analysis of a pair of periodically driven BECs in a double well potential. The periodic driving induces a synthetic gauge field in the Fock space of the system which is characterized by the boson occupation numbers of each well. For convenience we make the assumption that particle number is conserved, however, this is not necessary and our results are stable against some particle loss. Due to the particle number conservation the Fock states of each BEC are uniquely identified by a single number. These numbers serve as the coordinates in a two-dimensional (2D) FSL, akin to the site label of a 2D lattice in tight-binding models. We show that for low energies the driven system mimics that of a non-interacting ultracold atomic gas in a rotating harmonic trap (rotating traps were an early method used to create synthetic gauge fields in BECs [62, 63]). Within the rotating frame, the Coriolis force assumes the role of the Lorentz force experienced by a charged particle moving in a uniform magnetic field. However, this process has inherent limitations because the rotation exerts a centrifugal force on the atoms causing the atomic cloud to expand. As the rotation speed increases there comes a point called the centrifugal limit where the cloud becomes unstable and flies apart. In interacting ultracold gases vortex lattices can form and the centrifugal limit marks the point where the number of vortices is comparable to the number of particles in the gas. The gas becomes strongly correlated in this case which is referred to as the quantum Hall regime because states analogous to fractional quantum Hall states can form. Since our system resembles a non-interacting rotating gas, there is no vortex lattice present. However, when the FSL version of the centrifugal limit is passed, a transition is triggered and the ground state forms a vortex. The stability of the vortex state depends on higher order ‘trapping’ terms and finite size effects which forces the FSL to have hard wall edges. We quantify this transition in terms of the angular momentum in Fock space as well as the entanglement entropy between the two BECs with both showing a sudden increase in the vortex state.
@@
In a previous paper we predicted that vortices occur generically in the Fock space of spin systems following a sudden quench, the vortices being part of the fine structure of quantum caustics [61]. However, no attempt was made to specifically control the vortices. By contrast, here we suggest a scheme for generating tuneable artificial gauge potentials in Fock space that lead to controllable quantized vortices. This adds to the arsenal of quantum simulations that can be performed in artificial dimensions.
@@
The rest of this paper is laid out as follows: in Sec. II we describe the basic model that is assumed throughout this paper, namely two different but mutually interacting BECs in a driven double well potential. In Sec. III we start from the Floquet operator describing the periodically driven system and derive from it an effective Hamiltonian in Fock space for the total system that contains an artificial gauge potential. We compare various analytical predictions which can be made from this effective Hamiltonian against the exact Floquet theory in the main results section, Sec. IV. These results include the allowed energies and angular momenta of the vortex state, density and phase plots of the wavefunctions in Fock space, the dependence of the vortex transition on the interactions and number of particles and finally we show how the onset of a vortex in Fock space is associated with a jump in the entanglement entropy between the two BECs. We give our conclusions in Sec. V. There are also two appendices which give details of two of the calculations presented in the main body of the paper.
@@ II Model
The system we investigate consists of a pair of BECs in a double well potential undergoing time-periodic modulation. A single BEC in a static double well potential can be thought of as a bosonic version of the Josephson junction [64], and has been successfully realized experimentally by trapping atoms in either an actual double well potential [65, 66] or by using two different internal states of the atoms in a single well trap [67]. In our case we assume each of the two BECs is made from a different species of atom (alternative schemes involving a single species of atom in incoherent mixtures of two different internal states in a double well potential or even one species with four different internal states in a single well trap are conceivable, as long as no interconversion between the two ‘species’ takes place). For simplicity, we will assume that there are an equal number of $N$ particles of each species present, however this is not necessary and any difference will simply cause the FSL to be rectangular in shape rather than square.
@@
The modulation scheme we will use involves alternating between repulsive and attractive interactions between the two BECs, alongside periodic pulses controlling the tunneling of each BEC. The effect of one period $T$ of the modulation can be written in terms of the Floquet operator
@@
We take all of the parameters to be positive, so that $U>0$ controls the strength of always attractive intra-species interactions, whereas $W$ controls the inter-species interactions which alternate between attractive and repulsive. The tunneling between each well is controlled by $J$ and is pulsed once for each BEC each period for a duration $\delta t\ll 1$. We set both $U$ and $J$ to be the same for each BEC, again for convenience, but they are not required to be. However, it will become evident that there are important values of the ratio $U/J$ that cause the ground state wave function to change significantly, and so the ratio cannot be arbitrarily chosen.
@@
The Floquet operator can be broken down into five steps: 1) attractive intra- and inter-species interactions for a duration of $T/4$, 2) tunneling of one BEC for a duration of $\delta t$, 3) attractive intra- and repulsive inter-species interactions for a duration $T/2$, 4) tunneling of the other BEC for a duration of $\delta t$, 5) another attractive intra- and inter-species interactions for a duration of $T/4$. The combination of these steps results in a swirling effect in the 2D Fock space. This effect is reminiscent of the response of electrons moving in 2D in response to a uniform magnetic field pointing perpendicular to the plane. One of us (JM) has previously suggested a similar driving scheme to simulate large magnetic fields in the same 2D Fock space to investigate quantum Hall physics [48]. In this work, our main focus will instead be on simulating weak magnetic fields involving a small fraction of a flux quantum through each plaquette in the FSL.
@@
Achieving precise control over the parameters of the system is important for the physical implementation of the Floquet operator. Fortunately, the field of ultracold atoms has seen substantial progress in control techniques over the past couple of decades, and high levels of control can be accomplished. Species-specific trapping potentials [70] can be employed to independently control each BEC, while the pulse process can be realized by abruptly modifying the amplitude of the laser that forms the barrier between the wells for each BEC. Species-specific traps and periodic driving have been used in recent experiments involving ultracold atoms in double well potentials to simulate lattice gauge theories [71, 72]. Both the intra- and inter-species interactions can be controlled by using an applied magnetic field to alter the scattering length through a Feshbach resonance [73] where in the latter case an AC-magnetic field is needed to generate the periodic driving. This technique has previously been used to create ultracold molecules where the interaction energy is driven at the association frequency between atoms of the same species [74] and of two different species [75, 76].
III The effective Hamiltonian in Fock space
@@
As previously alluded to, the periodic driving is capable of producing a synthetic magnetic field in the 2D Fock space formed by the eigenstates of the $\hat{J}_{z}$ and $\hat{S}_{z}$ operators. Here, we will demonstrate the existence of the magnetic field explicitly by deriving a low energy effective Hamiltonian with a state dependent vector potential. To start, we note that the Floquet operator can be written in terms of an effective Hamiltonian, $\hat{U}_{F}=e^{-i\hat{H}_{\mathrm{eff}}\delta t}$. Unfortunately, the exact form of $\hat{H}_{\mathrm{eff}}$ contains an infinite number of terms which can be seen by combining all of the exponentials in Eq. (2) using the Baker-Campbell-Hausdorff formula [77]. However, for very short pulse duration, $\delta t$, the effective Hamiltonian can be well approximated by its zeroth order term in $\delta t$ giving (a derivation is provided in Appendix A)
@@
where the new scaled parameters are $u=\frac{UNT}{J\delta t}$ and $w=\frac{WNT}{4}$. We have explicitly written the Hamiltonian in terms of the Schwinger spin raising and lowering operators to highlight the phase factors which are analogous to Peierls factors arising in tight-binding models in the presence of a gauge potential. Indeed, it has been demonstrated that the terms within the square brackets of Eq. (LABEL:eq:Pham) closely resemble the Harper-Hofstadter model and the spectrum generated from them gives rise to the celebrated Hofstadter’s butterfly [48].
The synthetic magnetic flux per plaquette in the FSL can be calculated by finding the phase accumulated from one loop around a plaquette
@@
The synthetic magnetic field strength can be controlled by the driving period $T$, or by the inter-species interaction strength $W$, which also plays the role of the driving amplitude. We put $w=1$ for the remainder of the paper. As a consequence, owing to the presence of $N$ in $w$, this implies that the driving period is short or the inter-species interactions are weak, or both. This in turn means that the magnetic flux per plaquette becomes small, enabling us to perform a Taylor expansion of the phase factors in Eq. (LABEL:eq:Pham), while neglecting terms beyond quadratic order in $w/N$,
@@
The weak magnetic field allows us to further approximate Eq. (8) with the continuum approximation. This is because the magnetic length, $l_{B}=\sqrt{1/B}=\sqrt{N/2w}$ is much larger than the FSL lattice spacing, so eigenfunctions of the Floquet operator do not ‘see’ the discretization of the lattice. To make the continuum approximation, we first switch from the basis of the number difference between the two wells (i.e. the left/right modes) to the basis of the number difference between the even and odd modes (i.e.  the even and odd combinations of the left/right modes) which play the distinguished role of being the single particle ground and excited states in this two mode system. This is accomplished with the unitary transformation $e^{-i\frac{\pi}{2}\left(\hat{J}_{y}+\hat{S}_{y}\right)}\hat{H}_{\mathrm{eff}}e^{i\frac{\pi}{2}\left(\hat{J}_{y}+\hat{S}_{y}\right)}$. For the first species this results in the spin operator transformations $\hat{J}_{x}\to-\hat{J}_{z}^{\prime},\hat{J}_{y}\to\hat{J}_{y}^{\prime},\hat{J}_{z}\to\hat{J}_{x}^{\prime}$ where the prime indicates the operators are in the ground and excited state basis. Next, we perform the Holstein Primakoff transformation [78]
@@
which maps the spin operators to a single boson mode where $\hat{a}$ annihilates a boson in that mode. Assuming that we are dealing only with low energy states such that $\langle\hat{a}^{\dagger}\hat{a}\rangle/N\ll 1$, the new spin annihilation operator can be approximated as $\hat{J}_{-}^{\prime}\approx\sqrt{N}\hat{a}^{\dagger}$. The various components of the vector spin operator $\bm{\hat{J}^{\prime}}$ in the ground and excited state basis can likewise be approximated as $\hat{J}_{x}^{\prime}\approx\sqrt{N/2}\,\hat{x},\hat{J}_{y}^{\prime}\approx\sqrt{N/2}\,\hat{p}_{x}$ and $\hat{J}_{z}^{\prime}=-N/2+(\hat{p}_{x}^{2}+\hat{x}^{2}-1)/2$ where $\hat{x}=(\hat{a}^{\dagger}+\hat{a})/\sqrt{2}$ and $\hat{p}_{x}=i(\hat{a}^{\dagger}-\hat{a})/\sqrt{2}$ are the quadrature bosonic operators.
@@
where we have omitted constant terms and terms of $\mathcal{O}\left(1/N\right)$. The derivation of Eq. (10) from Eq. (LABEL:eq:Pham) can be thought of as analogous to going to the Landau level regime in the Harper-Hofstadter model with a harmonic trap. The position and momentum operators are $\bm{\hat{r}}=(\hat{x},\hat{y})$, $\bm{\hat{p}}=(\hat{p}_{x},\hat{p}_{y})$ and $\bm{\hat{A}}=w/2(-\hat{y},\hat{x})$ is the vector potential operator. For sufficiently large BECs, $\hat{H}_{\mathrm{eff}}$ is the Hamiltonian of a harmonically trapped charged particle in a uniform magnetic field, $B=\bm{\nabla}\times\bm{A}=w$ pointing in the direction perpendicular to the plane of the FSL. The missing factor of $2/N$ compared to the magnetic field calculated from Eq. (7) comes from the scaling of the FSL area by $2/N$ when transforming from the Schwinger spin operators to the quadrature operators.
The factor of unity in the harmonic trap strength arises due to the state-space being composed of many-body Fock states. Consequently, the tunneling between Fock states becomes state dependent, as indicated by the $C_{\pm}(m)$ factors in Eq. (1). This means that the tunneling becomes weaker as the edges are approached at $m_{\mathrm{edge}}=\pm N/2$ and this weakening produces the effective harmonic trap around $m=n=0$. The attractive intra-species interactions, characterized by strength $u$, also have harmonic dependence in Fock space but with the opposite sign and hence work against the confinement. This is because attractive intra-species interactions makes it energetically favorable for the particles in each BEC to clump into one well which promotes the state being closer to the edges.
@@
Nonetheless, there exists one major difference that will play a crucial role in the formation of a vortex in the ground state of the system. In rotating gas experiments it is often the goal to have the rotation frequency approach the transverse trapping frequency, $\Omega\to\omega_{\bot}$, in order to simulate Landau level physics. The point where $\Omega=\omega_{\bot}$ is the centrifugal limit and cannot be reached because the system becomes unstable there. In $\hat{H}_{\mathrm{eff}}$, the point $u=1$ is analogous to the centrifugal limit. In the thermodynamic limit $N\rightarrow\infty$ it marks the point where a ground state quantum phase transition takes place from a normal phase ($u<1$) where $\langle\hat{S}_{z}\rangle$, $\langle\hat{J}_{z}\rangle=0$ to a symmetry broken phase ($u>1$) where $|\langle\hat{S}_{z}\rangle|$, $|\langle\hat{J}_{z}\rangle|\neq 0$. However, for finite $N$ there are higher-order terms which we neglected in the derivation of Eq. (10) which stabilize the ground state. Indeed, finite system sizes must stabilize the system because the FSL has hard wall edges since one cannot have more particles in a well than exist in the BEC. Therefore, the effective full ‘trapping potential’ in the FSL for the finite $N$ case more closely resembles the combination of a quadratic-plus-quartic potential [79, 80] and a box trap [81].

### v2

@@ Abstract
We perform a theoretical study of the coupled dynamics of two species of Bose-Einstein condensates (BECs) in a double-well potential where both the tunneling and the interatomic interactions are driven periodically in time. The population difference between the wells of each species gives rise to a two-dimensional lattice in Fock space with dimensions given by the number of atoms in each BEC. We use a Floquet analysis to derive an effective Hamiltonian that acts in this Fock space and find that it contains an artificial gauge field. This system simulates noninteracting particles in a tight-binding lattice subject to an additional harmonic potential and vector potential. When the intra-species interactions are attractive there is a critical value at which the ground state undergoes a transition from a Gaussian state to a quantized vortex state in Fock space. The transition can be quantified in terms of the angular momentum and the entanglement entropy of the ground state with both showing sudden jumps as the intra-species interactions become stronger. The stability of the vortex state vanishes in the thermodynamic limit.
@@ I Introduction
Another useful tool for quantum simulation is provided by artificial gauge potentials. These mimic the action of electric and magnetic fields on charges irrespective of whether the underlying physical particles are charged or not [20, 21, 22, 23]. For ultracold atoms, a popular experimental configuration is to load them into an optical lattice which can then be periodically modulated to generate effective gauge potentials [24, 25, 26, 27, 28, 29, 30, 31, 32]. The ability to simulate gauge fields in a controlled and tuneable manner has led to new proposals for investigating topological phases [33, 34, 35, 36, 37] and exotic phases of matter [38, 39], as well as experiments studying quantum Hall physics on entirely different physical platforms to their original settings [11, 40]. Artificial gauge fields not only provide a versatile experimental method for probing fundamental aspects of quantum mechanics and condensed matter physics but also enable new technological applications in quantum information processing and quantum simulation [41, 42, 43].
@@
Fock states provide yet another possible set of states that can be used as building blocks for synthetic dimensions. Unlike the more traditionally used internal and external states listed above, Fock states are defined via the occupation numbers of particles in the different available modes leading to the concept of a Fock-state lattice (FSL) [45, 44]. Systems with a large number of particles and a limited number of modes are particularly well-suited for constructing large synthetic dimensions in this way because each mode acts as a single synthetic dimension, and the number of particles that can occupy a mode serves as the length of the dimension. In this context, particles with bosonic statistics are preferable due to their ability to occupy the same mode simultaneously. One example of a FSL is provided by photon states in optical cavities; when two or three optical cavities are coupled via a two-level atom such a FSL can realize the Su-Schrieffer-Heeger (SSH) model or a Lifshitz topological phase transition between a semimetal and a three band insulator, respectively [45, 46]. Likewise, experiments with superconducting circuits have successfully constructed one- and two-dimensional (2D) FSLs which can mimic both the SSH and Haldane models [47]. Of particular relevance to the current paper is a recent proposal by one of us (JM) based on two species of atomic Bose-Einstein condensates (BECs) confined in a double-well potential which realizes a two-dimensional FSL [48]. This system can simulate strong synthetic gauge fields by intense periodic driving of the interactions between the two species of atoms leading to topological behavior in the form of chiral edge states.
[…]
@@
In this paper we study vortices in a FSL by combining elements of all the above-mentioned systems. Specifically, we perform a theoretical analysis of a pair of periodically driven BECs in a double-well potential. The periodic driving induces a synthetic gauge field in the Fock space of the system which is characterized by the boson occupation numbers of each well. For convenience we make the assumption that particle number is conserved, however, this is not necessary and our results are stable against some particle loss. Due to the particle number conservation the Fock states of each BEC are uniquely identified by a single number. These numbers serve as the coordinates in a 2D FSL, akin to the site label of a 2D lattice in tight-binding models. We show that for low energies the driven system mimics that of a non-interacting ultracold atomic gas in a rotating harmonic trap (rotating traps were an early method used to create synthetic gauge fields in BECs [61, 62]). Within the rotating frame, the Coriolis force assumes the role of the Lorentz force experienced by a charged particle moving in a uniform magnetic field. However, this process has inherent limitations because the rotation exerts a centrifugal force on the atoms causing the atomic cloud to expand. As the rotation speed increases there comes a point called the centrifugal limit where the cloud becomes unstable and flies apart. In interacting ultracold gases vortex lattices can form and the centrifugal limit marks the point where the number of vortices is comparable to the number of particles in the gas. The gas becomes strongly correlated in this case, which is referred to as the quantum Hall regime because states analogous to fractional quantum Hall states can form. Since our system resembles a non-interacting rotating gas, there is no vortex lattice present. However, when the FSL version of the centrifugal limit is passed, a transition is triggered and the ground state forms a vortex. The stability of the vortex state depends on higher-order ‘trapping’ terms and finite-size effects, which forces the FSL to have hard wall edges. We quantify this transition in terms of the angular momentum in Fock space as well as the entanglement entropy between the two BECs with both showing a sudden increase in the vortex state.
@@
In a previous paper we predicted that vortices can occur in the Fock space of spin systems following a sudden quench, the vortices being part of the fine structure of quantum caustics [63]. However, no attempt was made to specifically control the vortices. By contrast, here we suggest a scheme for generating tuneable artificial gauge potentials in Fock space that lead to controllable quantized vortices. This adds to the arsenal of quantum simulations that can be performed in artificial dimensions.
@@
The rest of this paper is laid out as follows: in Sec. II we describe the basic model that is assumed throughout this paper, namely two different but mutually interacting BECs in a driven double-well potential. In Sec. III we start from the Floquet operator describing the periodically driven system and derive from it an effective Hamiltonian in Fock space for the total system that contains an artificial gauge potential. We compare various analytical predictions which can be made from this effective Hamiltonian against the exact Floquet theory in the main results section, Sec. IV. These results include the allowed energies and angular momenta of the vortex state, density and phase plots of the wave functions in Fock space and the dependence of the vortex transition on the interactions and number of particles, and finally we show how the onset of a vortex in Fock space is associated with a jump in the entanglement entropy between the two BECs. We give our conclusions in Sec. V. There are also two appendices which give details of two of the calculations presented in the main body of the paper.
@@ II Model
The system we investigate consists of a pair of BECs in a double-well potential undergoing time-periodic modulation. A single BEC in a static double-well potential can be thought of as a bosonic version of the Josephson junction [64], and has been successfully realized experimentally either by trapping atoms in an actual double-well potential [65, 66] or by using two different internal states of the atoms in a single-well trap [67]. In our case we assume each of the two BECs is made from a different species of atom (alternative schemes involving a single species of atom in incoherent mixtures of two different internal states in a double-well potential or even one species with four different internal states in a single-well trap are conceivable, as long as no interconversion between the two ‘species’ takes place). For simplicity, we will assume that there are an equal number of $N$ particles of each species present, however this is not necessary and any difference will simply cause the FSL to be rectangular in shape rather than square.
@@
The modulation scheme we will use involves alternating between repulsive and attractive interactions between the two BECs, alongside periodic pulses controlling the tunneling of each BEC. If the tunneling interval $\delta t$ is very short such that no other dynamics takes places over the interval, then the effect of one period of the modulation can be written in terms of the Floquet operator
@@
We take all of the parameters to be positive, so that $U>0$ controls the strength of attractive intra-species interactions, whereas $W$ controls the inter-species interactions which alternate between attractive and repulsive. The tunneling between each well is controlled by $J$ and is pulsed once for each BEC each period for a duration $\delta t$. We set both $U$ and $J$ to be the same for each BEC, again for convenience, but they are not required to be. However, it will become evident that there are important values of the ratio $U/J$ that cause the ground-state wave function to change significantly, and so the ratio cannot be arbitrarily chosen.
@@
The Floquet operator can be broken down into five steps: 1) attractive intra- and inter-species interactions for a duration of $T/4$, 2) tunneling of one BEC for a duration of $\delta t$, 3) attractive intra- and repulsive inter-species interactions for a duration $T/2$, 4) tunneling of the other BEC for a duration of $\delta t$, 5) other attractive intra- and inter-species interactions for a duration of $T/4$. The combination of these steps results in a swirling effect in the 2D Fock space. This effect is reminiscent of the response of electrons moving in two dimensions in response to a uniform magnetic field pointing perpendicular to the plane. One of us (JM) has previously suggested a similar driving scheme to simulate large magnetic fields in the same 2D Fock space to investigate quantum Hall physics [48]. In this paper, our main focus will instead be on simulating weak magnetic fields involving a small fraction of a flux quantum through each plaquette in the FSL.
@@
Achieving precise control over the parameters of the system is important for the physical implementation of the Floquet operator. Fortunately, the field of ultracold atoms has seen substantial progress in control techniques over the past couple of decades, and high levels of control can be accomplished. For example, Floquet engineering of the tunneling between the wells can be achieved by modulating the lasers creating a barrier [70], and species-specific traps and periodic driving have been used in experiments involving ultracold atoms in double-well potentials to simulate lattice gauge theories in a setup which is quite similar to that which we are proposing here [71] (see also the review [72]). Ultracold atoms also offer the ability to control interactions between being repulsive and attractive by using a magnetic field to tune the interatomic scattering length through a Feshbach resonance. In our case we need to separately control the intra- and inter-species interactions such that the intra-species interactions are always attractive but the inter-species interactions periodically alternate in sign. This can achieved by choosing two species that each have attractive interactions within a certain range of magnetic-field magnitude, but there is an inter-species Feshbach resonance that lies within this range. For example, certain states of 133Cs, 7Li, and 39K have considerable ranges of attractive interactions and also interspecies Feshbach resonances [73, 74, 75]. Time-dependent control of scattering lengths has also been demonstrated, e.g.  a time-modulated magnetic field has been used to generate periodic driving of the interactions [76, 77]. These types of techniques were originally used to create ultracold molecules where the interaction energy is driven at the association frequency between atoms of the same species [78]. More recently, and of great pertinence to the present proposal, they have also been used to control the inter-species interactions in dual species BECs of 87Rb and 41K [79], of 87Rb and 85Rb [80], of 168Yb and 174Yb [81], of 87Rb and 133Cs [82], of 23Na and 87Rb [83], of 87Rb and 84Sr and 87Rb and 88Sr [84], of 39K and 87Rb [85], and of 39K and 41K [86].
In order to observe vortices in Fock space one should be able to measure the probability distribution as a function of the coordinates $m$ and $n$ which are the number differences between the two wells for the two species. For a single species this has previously been achieved by using absorption imaging [65] and also resonant fluorescence detection [87]. In the latter case, resolution at the level of a single atom number was reported in a sample of 1200 atoms. To apply this technique to a dual species condensate it is sufficient that the two species have transitions which differ enough in frequency that they can be individually addressed by resonant lasers (which is often the case). It should be remembered that a single measurement of the number difference for each species will yield a single number: in order to build up a probability distribution these measurements must be repeated many times, each time on a system which has been prepared as identically as possible to all the other measurements.
III Effective Hamiltonian in Fock space
@@
The periodic driving is capable of producing a synthetic magnetic field in the 2D Fock space formed by the eigenstates of the $\hat{J}_{z}$ and $\hat{S}_{z}$ operators. Here, we will demonstrate the existence of the magnetic field explicitly by deriving a low-energy effective Hamiltonian with a state dependent vector potential. To start, we note that the Floquet operator can be written in terms of an effective Hamiltonian, $\hat{U}_{F}=e^{-i\hat{H}_{\mathrm{eff}}\delta t}$. Unfortunately, the exact form of $\hat{H}_{\mathrm{eff}}$ contains an infinite number of terms which can be seen by combining all of the exponentials in Eq. (2) using the Baker-Campbell-Hausdorff formula [88]. However, for very short pulse duration, $\delta t$, the effective Hamiltonian can be well approximated by its zeroth-order term in $\delta t$, giving (a derivation is provided in Appendix A)
@@
where the new scaled parameters are $u=\frac{UNT}{J\delta t}$ and $w=\frac{WNT}{4}$. We have explicitly written the Hamiltonian in terms of the Schwinger spin raising and lowering operators to highlight the phase factors which are analogous to Peierls factors arising in tight-binding models in the presence of a gauge potential. Indeed, it has been demonstrated that the terms within the square brackets of Eq. (LABEL:eq:Pham) closely resemble the Harper-Hofstadter model and the spectrum generated from them gives rise to the celebrated Hofstadter butterfly [48].
The synthetic magnetic flux per plaquette in the FSL can be calculated by finding the phase accumulated from one loop around a plaquette:
@@
The synthetic magnetic-field strength can be controlled by the driving period $T$, or by the inter-species interaction strength $W$, which also plays the role of the driving amplitude. We set $w=1$ for the remainder of the paper. As a consequence, owing to the presence of $N$ in $w$, this implies that the driving period is short or the inter-species interactions are weak, or both. This in turn means that the magnetic flux per plaquette becomes small, enabling us to perform a Taylor expansion of the phase factors in Eq. (LABEL:eq:Pham), while neglecting terms beyond quadratic order in $w/N$:
@@
The weak magnetic field allows us to further approximate Eq. (8) with the continuum approximation. This is because the magnetic length, $l_{B}=\sqrt{1/B}=\sqrt{N/2w}$ is much larger than the FSL lattice spacing, so eigenfunctions of the Floquet operator do not ‘see’ the discretization of the lattice. To make the continuum approximation, we first switch from the basis of the number difference between the two wells (i.e. the left and right modes) to the basis of the number difference between the even and odd modes (i.e.  the even and odd combinations of the left and right modes) which play the distinguished role of being the single-particle ground and excited states in this two mode system. This is accomplished with the unitary transformation $e^{-i\frac{\pi}{2}\left(\hat{J}_{y}+\hat{S}_{y}\right)}\hat{H}_{\mathrm{eff}}e^{i\frac{\pi}{2}\left(\hat{J}_{y}+\hat{S}_{y}\right)}$. For the first species this results in the spin operator transformations $\hat{J}_{x}\to-\hat{J}_{z}^{\prime},\hat{J}_{y}\to\hat{J}_{y}^{\prime},\hat{J}_{z}\to\hat{J}_{x}^{\prime}$ where the prime indicates the operators are in the ground and excited-state basis. Next, we perform the Holstein-Primakoff transformation [89]
@@
which maps the spin operators to a single boson mode where $\hat{a}$ annihilates a boson in that mode. Assuming that we are dealing only with low-energy states such that $\langle\hat{a}^{\dagger}\hat{a}\rangle/N\ll 1$, the new spin annihilation operator can be approximated as $\hat{J}_{-}^{\prime}\approx\sqrt{N}\hat{a}^{\dagger}$. The various components of the vector spin operator $\bm{\hat{J}^{\prime}}$ in the ground and excited state basis can likewise be approximated as $\hat{J}_{x}^{\prime}\approx\sqrt{N/2}\,\hat{x},\hat{J}_{y}^{\prime}\approx\sqrt{N/2}\,\hat{p}_{x}$ and $\hat{J}_{z}^{\prime}=-N/2+(\hat{p}_{x}^{2}+\hat{x}^{2}-1)/2$ where $\hat{x}=(\hat{a}^{\dagger}+\hat{a})/\sqrt{2}$ and $\hat{p}_{x}=i(\hat{a}^{\dagger}-\hat{a})/\sqrt{2}$ are the quadrature bosonic operators.
@@
where we have omitted constant terms and terms of $\mathcal{O}\left(1/N\right)$. The derivation of Eq. (10) from Eq. (LABEL:eq:Pham) is analogous to going to the Landau-level regime in the Harper-Hofstadter model with a harmonic trap. The position and momentum operators are $\bm{\hat{r}}=(\hat{x},\hat{y})$ and $\bm{\hat{p}}=(\hat{p}_{x},\hat{p}_{y})$, and $\bm{\hat{A}}=w/2(-\hat{y},\hat{x})$ is the vector potential operator. For sufficiently large BECs, $\hat{H}_{\mathrm{eff}}$ is the Hamiltonian of a harmonically trapped charged particle in a uniform magnetic field, $B=\bm{\nabla}\times\bm{A}=w$ pointing in the direction perpendicular to the plane of the FSL. The missing factor of $2/N$ compared to the magnetic field calculated from Eq. (7) comes from the scaling of the FSL area by $2/N$ when transforming from the Schwinger spin operators to the quadrature operators.
The factor of unity in the harmonic trap strength arises due to the state space being composed of many-body Fock states. Consequently, the tunneling between Fock states becomes state dependent, as indicated by the $C_{\pm}(m)$ factors in Eq. (1). This means that the tunneling becomes weaker as the edges are approached at $m_{\mathrm{edge}},n_{\mathrm{edge}}=\pm N/2$ and this weakening produces the effective harmonic trap around $m=n=0$. The attractive intra-species interactions, characterized by strength $u$, also have harmonic dependence in Fock space but with the opposite sign and hence work against the confinement. This is because attractive intra-species interactions make it energetically favorable for the particles in each BEC to clump into one well and these states are closer to the edges of the FSL.
@@
Nonetheless, there exists one major difference that will play a crucial role in the formation of a vortex in the ground state of the system. In rotating gas experiments it is often the goal to have the rotation frequency approach the transverse trapping frequency, $\Omega\to\omega_{\bot}$, in order to simulate Landau level physics. The point where $\Omega=\omega_{\bot}$ is the centrifugal limit and cannot be reached because the system becomes unstable there. In $\hat{H}_{\mathrm{eff}}$, the point $u=1$ is analogous to the centrifugal limit. In the thermodynamic limit $N\rightarrow\infty$ it marks the point where a ground-state quantum phase transition takes place from a normal phase ($u<1$) where $\langle\hat{S}_{z}\rangle$, $\langle\hat{J}_{z}\rangle=0$ to a symmetry broken phase ($u>1$) where $|\langle\hat{S}_{z}\rangle|$, $|\langle\hat{J}_{z}\rangle|\neq 0$. However, for finite $N$ there are higher-order terms which we neglected in the derivation of Eq. (10) which stabilize the ground state. Indeed, finite system sizes must stabilize the system because the FSL has hard wall edges since one cannot have more particles in a well than exist in the BEC. Therefore, the effective full ‘trapping potential’ in the FSL for the finite $N$ case more closely resembles the combination of a quadratic-plus-quartic potential [90, 91] and a box trap [92].